if 507 g FeCl2 were used up in the reaction FeCl2 + 2NaOH = Fe(OH)2(s) + 2NaCl, how many grams of NaCl would be made?
1
Expert's answer
2019-05-20T08:23:59-0400

507 g FeCl2 it is 507/(56+35.5*2) = 3.76mol. Based on the reaction equation of sodium chloride is formed 3,76*2=7,52mol, it is 7.52*(23+35.5) = 440g